OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -1 and a y-intercept of 2?

OpenStudy (anonymous):

I dont understand how to do it the options are y=2x-2 y=-2x+2 y=2x+2 If someone could show me how its done thatd be great

OpenStudy (triciaal):

x-intercept means when y = 0 y-intercept means when x = 0

OpenStudy (anonymous):

oh my goodness thank you so much that statement alone just helped me a ton

OpenStudy (triciaal):

can you find your 2 points on this line given the intercepts?

OpenStudy (anonymous):

so like 0=2(-1)-2, would that be the proper way to solve this?

OpenStudy (triciaal):

you have the points find the slope you know the slope is the steepness of the line and it is the same for any 2 points on that line slope,m = change in y /CORRESPONDING change in x for 2 points (x1, y1) and (x2, xy2) the slope is (y2 - y1)/(x2-x1) then with this value for m and any point on the line (x, y) so can write the equation of the line point slope form (y-y1) = m(x-x1) note that you could use (x2, y2) for the equation you pick a point

OpenStudy (anonymous):

x/x-intercept+y/y-intercept =1 is the general formula, from herer putting the values of xcintercept and y inercept you wiull get the result

OpenStudy (triciaal):

what are your 2 points given the intercepts?

OpenStudy (triciaal):

What is the equation of the line with an x-intercept of -1 and a y-intercept of 2?

OpenStudy (anonymous):

x/(-1)+y/2=1 implies 2x+2=y is the result

OpenStudy (anonymous):

@Jonnyloo

OpenStudy (triciaal):

@Jonnyloo do you understand the steps outlined above? do you have any questions?

OpenStudy (triciaal):

x-intercept of -1 point (-1, 0) y-intercept of 2 point (0, 2) slope (2 - 0)/ (0--1) = 2/1 = 2 using point (0, 2) (y - 2) = 2(x-0) equation y - 2 = 2 x in the form y = mx + b gives y = 2x +2 = choice 3